OpenStudy (anonymous):

5(t + 1) + 8 = 3(t − 2) + 6 please show work. Solve and check. Label any contradictions or identities

OpenStudy (wach):

Try distributing 5 into (t + 1) and 3 into (t - 2).

OpenStudy (anonymous):

5(t+1)+8= 3(t-2)+6 5t + 5 + 8 = 3t - 6 + 6 5t + 13 = 3t 13 = 3t - 5t 13 = -2t t = 13/-2 t = -6.5 this is what i came up with i was just checking to see if it was right

OpenStudy (wach):

Nice job! I believe it is correct :)

OpenStudy (anonymous):

thank you